OpenStudy (anonymous):

Solve the quadratic equation 3x2 + 5x + 2 = 0 by factoring.

OpenStudy (anonymous):

3x^2 + 5x + 2 = 0 (3x + 2) (x + 1) = 0

OpenStudy (anonymous):

Try to go from here, by setting each term equal to 0 separately.